SOLUTION: C and Y candy company mixes candy that costs $6.00 per kg with candy that costs $4.50 per kg. How many kg's of each are needed to make a 3kg box that costs $15.00 ?
you need to do it in a chart

Answer by lwsshak3(11628)

SOLUTION: C and Y candy company mixes candy that costs $6.00 per kg with candy that costs $4.50 per kg. How many kg's of each are needed to make a 3kg box that costs $15.00 ?
you need to do it in a chart
**
.....$/kg.......kg..........$cost
.....6.00.........x............6x
.....4.50........3-x.........4.5(3-x)
.................... 3.............15
6x+4.5(3-x)=15
6x+13.5-4.5x=15
1.5x=1.5
x=1
3-x=2
kg of $6 candy needed=1
kg of $4.50 candy needed=2